# Kleene Algebrae, Kleene Modules, and Morita Equivalence

### v1

@@ 2 Kleene modules
Fix Kleene algebrae $A$, $B$, $C$. A tensor product is a functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$ such that for each $N\in{{}_{B}\text{Mod}_{C}}$, the functors $\lambda M.\,F(M,N)$ and $\lambda P.\,\operatorname{Hom}(N,P)$ are adjoint.
@@
Note that it follows from the universal property for adjoints that tensor products, if they exist, are unique up to isomorphism. One can also define tensor products producing left and right Kleene modules, with types ${{}_{A}\text{Mod}_{B}}\times\text{Mod}_{B}\rightarrow\text{Mod}_{A}$ and ${{}_{A}\text{Mod}}\times{{}_{A}\text{Mod}_{B}}\rightarrow{{}_{A}\text{Mod}}$, in the obvious way. We omit the definition since it won’t be used in the sequel.
@@
Defining tensor products of Kleene modules doesn’t guarantee they exist; this is established next.
@@
For any Kleene algebrae $A$, $B$, $C$ there is a tensor product functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$.
@@
We proceed by direct construction. Fix $M\in{{}_{A}\text{Mod}_{B}}$ and $N\in{{}_{B}\text{Mod}_{C}}$. Let $\Xi$ be the Kleene module freely generated by $M\times N$, and let $\Phi$ be the congruence generated by the relations
$(m+m^{\prime},n)\Phi(m,n)+(m^{\prime},n),(m,n+n^{\prime})\Phi(m,n)+(m,n^{\prime}),(mb,n)\Phi(m,bn)$
for $m,m^{\prime}\in M$, $n,n^{\prime}\in N$, and $b\in B$. Define $M\otimes_{B}N=\Xi/\Phi$.
@@
$\operatorname{Hom}(M\otimes_{B}N,P)$$\operatorname{Hom}(M,\operatorname{Hom}(N,P))$$\operatorname{Hom}(M^{\prime}\otimes_{B}N,P^{\prime})$$\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$$\scriptstyle{\operatorname{Hom}(\alpha\otimes\text{id},\alpha)}$$\scriptstyle{\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))}$
@@
$=\beta(\varphi(\alpha(m^{\prime})\otimes n))$
@@
It remains to find an inverse to the natural transformation $\varphi\mapsto\overline{\varphi}$. The natural candidate is given for $M\in{{}_{A}\text{Mod}_{B}}$, $P\in{{}_{B}\text{Mod}_{C}}$, $\psi\in\operatorname{Hom}(M,\operatorname{Hom}(N,P))$, $m\in M$, and $n\in N$ by $\hat{\psi}(m\otimes n)=\psi(m)(n)$. It is immediate that the transformation $\psi\mapsto\hat{\psi}$ is inverse to the transformation $\phi\mapsto\overline{\phi}$. It remains to verify that $\psi\mapsto\hat{\psi}$ is a natural transformation. For this we need to check that the following diagram commutes, where $M,M^{\prime}\in{{}_{A}\text{Mod}_{B}}$, $P,P^{\prime}\in{{}_{B}\text{Mod}_{C}}$, $\alpha\mathrel{:}M\rightarrow M^{\prime}$, $\beta\mathrel{:}P^{\prime}\rightarrow P$, and the horizontal arrows are given by application of the transformation $\psi\mapsto\hat{\psi}$.
@@
$\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))\hat{\phantom{|}}(\psi)(m\otimes n)$ $=\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))(\psi)(m)(n)$
@@
2.1 The Morita Category
Morita equivalence is of fundamental importance to ring theory, and based on the suspicion that it may be useful in the study of Kleene algebrae as well we develop it here. Morita equivalence for rings is obtained by replacing the morphisms of the category of rings by bimodules, and considering the induced notion of isomorphism. More specifically, the morphisms from a ring $R$ to a ring $S$ are the $(R,S)$-bimodules, and composition of bimodules is given by tensor product. For more on Morita equivalence of rings see [3].
The case of Morita equivalence for Kleene algebrae is a straightforward adaptation of that for rings.
Definition 4.
The Morita category of Kleene algebrae has objects Kleene algebrae, and for $A$ and $B$ Kleene algebrae the morphisms between them are all Kleene $(A,B)$-bimodules. The identity at an algebra $K$ is ${{}_{K}}K_{K}$, and composition of morphisms is given by tensor product.
To verify that this is indeed a category, we must check that tensor product is associative and has the bimodules ${{}_{K}}K_{K}$ as identities.
@@ 2.1 The Morita Category
Let $A$, $B$, $C$, $D$ be Kleene algebrae, and ${{}_{A}}M_{B}$, ${{}_{B}}N_{C}$, ${{}_{C}}P_{D}$ be Kleene bimodules. Then
@@
By the defining property of tensor products, for any Kleene module ${{}_{D}}Q_{E}$ we natural isomorphisms
$\operatorname{Hom}(({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D},{{}_{D}}Q_{E})\cong\operatorname{Hom}(M,\operatorname{Hom}(N,\operatorname{Hom}(P,Q)))$
and
$\operatorname{Hom}({{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D}),{{}_{D}}Q_{E})\cong\operatorname{Hom}(M,\operatorname{Hom}(N,\operatorname{Hom}(P,Q))).$
Hence $({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D}\cong{{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D})$. Also for any $a\in A$ and $m\in M$, $a\otimes m=a(1\otimes m)$, so the map $m\mapsto 1\otimes m$ is an isomorphism from $M$ to $A\otimes M$. ∎
As we have defined it, the Morita “category” isn’t actually a category, because the associativity and identity laws only hold up to isomorphism. In practice this won’t matter, since isomorphic Kleene algebrae and modules are regarded as “the same.” If one insists on working with a bona fide category, one may restrict to a skeleton of the Morita category, which includes precisely one element of each isomorphism class of objects and morphisms from the Morita category. Also, we said that the Morita category extends the usual category of Kleene algebrae, and must check that each homomorphism between Kleene algebrae gives rise to a Kleene module in a manner which sends a composition of homomorphisms to the tensor product of the corresponding modules.
@@
Let $A$, $B$ be Kleene algebrae and $h\mathrel{:}A\rightarrow B$ be a homomorphism. The homomorphism module $E_{h}$ is a Kleene $(A,B)$-bimodule whose elements and addition are given by the additive reduct of $B$, with the following multiplication actions on $b\in B$: For $a\in A$, $ab=h(a)b$. For $^{\prime}\in B, bb$$b^{\prime}=bb^{\prime}$ (just perform the multiplication in $B$). It is immediate that this givs $E_{h}$ the structure of a Kleene bimodule.
@@
Let $A$, $B$, $C$ be Kleene algebrae and $f\mathrel{:}A\rightarrow B$ and $g\mathrel{:}B\rightarrow C$ be homomorphisms. Then $E_{g\circ f}\cong E_{f}\otimes E_{g}$.
@@
$\phi(ac)$ $=1\otimes(g\circ f)(a)c$
@@
Now consider $b\otimes c\in E_{g}\otimes E_{f}$. Define $\psi\mathrel{:}E_{f}\otimes E_{g}\rightarrow E_{g\circ f}$ by $\psi(b\otimes c)=bc$, and note that this is well-defined because $b\otimes c=1\otimes bc$ for $b\in B$, $c\in C$. It is clearly additive and commutes with the right action of $C$, so we demonstrate that it commutes with the left action of $A$. Let $a\in A$ and note that
@@
$=\psi(f(a)\otimes bc)$
$=f(a)bc$
$=a\psi(b\otimes c)$
@@
$\phi(\psi(b\otimes c))=\phi(bc)=1\otimes bc=b\otimes c$
@@
Another way of expressing Morita equivalence for rings is as equivalence of the corresponding module categories. This carries over to Kleene algebrae, as we now argue. If $K$ is Morita-equivalent with $S$ via bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$, then for any left $K$-module ${{}_{K}}P$,
@@
As in the classical case, a Kleene algebra is Morita-equivalent with each of its finite-dimensional matrix rings. This can be viewed as explaining why the behaviour of finite automata over a Kleene algebrae, which can be defined by finite-dimensional square matrices, is fully expressible in terms of the base Kleene algebra.
@@
The $(M,K)$-bimodule $K^{n}$ and its dual $K^{n\circ}$ will witness this equivalence. First consider $K^{n}\otimes K^{n\circ}$. For $1\leq i\leq n$ let $e_{i}$ be the tuple which is zero everywhere except position $i$, where it is $1$. Note that the set $\{e_{i}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n}$. Let $e_{i}^{\circ}$ be the dual of this module element, that is, $e_{i}^{\circ}(e_{i})=1$ and $e_{i}^{\circ}(e_{j})=0$ for $j\neq i$. It is easy to show that the set $\{e_{i}^{\circ}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n\circ}$. For $1\leq i,j\leq n$, let $E_{ij}\in M$ be the matrix which is zero everywhere except at position $(i,j)$, where it is $1$. Note that these freely generate $M$, and that the set $\{e_{i}\otimes e_{j}^{\circ}\mathrel{:}1\leq i,j\leq n\}$ freely generates $K^{n}\otimes K^{n\circ}$, because only elements of $K$ can be transported across the tensor product. Define a map $\phi\mathrel{:}K^{n}\otimes K^{n\circ}$ by $\phi(e_{i}\otimes e_{j}^{\circ})=E_{ij}$; we show that $\phi$ is an isomorphism. Its inverse, naturally, will be given by $\psi(E_{ij})=e_{i}\otimes e_{j}^{\circ}$. By free generation $\phi$ and $\psi$ are well-defined homomorphisms, and it is immediate that they are inverses. Consequently $K^{n}\otimes K^{n\circ}\cong M$, which is half of a Morita equivalence.
For the other half we look at $K^{n\circ}\otimes K^{n}$, and our goal is to show this is isomorphic to $K$. For $e_{i}^{\circ}\otimes e_{j}\in K^{n\circ}\otimes K^{n}$, define $\alpha(e_{i}^{\circ}\otimes e_{j})$ to be 0 if $i\neq j$ and $1$ if $i=j$. By free generation this determines a homomorphism $\alpha\mathrel{:}K^{n\circ}\otimes K^{n}\rightarrow K$. Define also $\beta\mathrel{:}K\rightarrow K^{n\circ}\otimes K^{n}$ by $\beta(a)=a(e_{1}^{\circ}\otimes e_{1})$ for $a\in A$, which is well-defined by free generation. To complete the proof we show that $\alpha$ and $\beta$ are inverses. Certainly for $a\in A$ we have that
@@
To see that $\beta\circ\alpha=\text{id}$ it suffices to show that for any $i\leq n$, $e_{i}^{\circ}\otimes e_{i}=e_{1}^{\circ}\otimes e_{1}$. For this we show that
@@
$\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$ $=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
@@
$e_{i}^{\circ}\otimes e_{i}$ $=e_{i}^{\circ}\overline{e}_{i}\otimes e_{i}$
$=e_{i}^{\circ}\otimes\overline{e}_{i}e_{i}$
$=e_{i}^{\circ}\otimes\sum_{i=1}^{n}e_{i}$
$=e_{i}^{\circ}\otimes\overline{1}\sum_{i=1}^{n}e_{i}$
$=e_{i}^{\circ}\overline{1}\otimes\sum_{i=1}^{n}e_{i}$
$=\left(\sum_{i=1}^{n}e_{i}^{\circ}\right)\otimes\sum_{i=1}^{n}e_{i}.$
@@
There is a well-known characterization of Morita equivalence for rings which extends to semirings [3], [2]. To state it we need the following definition:
@@
Let $R$ be a semiring and $e\in R$ a multiplicative idempotent (so $e^{2}=e$). The idempotent $e$ is full if and only if $ReR=R$.

### v2

@@ 3.2 Tensor products of Kleene modules
Fix Kleene algebras $A$, $B$, $C$. A tensor product is a functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$ such that for each $N\in{{}_{B}\text{Mod}_{C}}$, the functors $M.\mapsto F(M,N)$ and $P\mapsto\operatorname{Hom}(N,P)$ form the left and right parts of an adjunction, respectively.
@@
Note that it follows from the universal property for adjoints that tensor products, if they exist, are unique up to isomorphism [16, IV.1]. One can also define tensor products producing left and right Kleene modules, with types ${{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}}\rightarrow{{}_{A}\text{Mod}}$ and $\text{Mod}_{A}\times{{}_{A}\text{Mod}_{B}}\rightarrow\text{Mod}_{A}$, in the obvious way. We omit the definition since it will not be used in the sequel.
@@
Defining tensor products of Kleene modules via an adjunction requirement does not guarantee they exist; this is established next.
@@
For any Kleene algebras $A$, $B$, $C$ there is a tensor product functor $F\mathrel{:}{{}_{A}\text{Mod}_{B}}\times{{}_{B}\text{Mod}_{C}}\rightarrow{{}_{A}\text{Mod}_{C}}$.
@@
We proceed by direct construction. Fix $M\in{{}_{A}\text{Mod}_{B}}$ and $N\in{{}_{B}\text{Mod}_{C}}$. Let $\Xi$ be the Kleene $(A,C)$-bimodule freely generated by $M\times N$, and let $\Phi$ be the congruence generated by the relations
$(m+m^{\prime},n)\mathrel{\Phi}(m,n)+(m^{\prime},n),(m,n+n^{\prime})\mathrel{\Phi}(m,n)+(m,n^{\prime}),(mb,n)\mathrel{\Phi}(m,bn)$
for $m,m^{\prime}\in M$, $n,n^{\prime}\in N$, and $b\in B$. Define $M\otimes_{B}N=\Xi/\Phi$. This immediately has the structure of a semiring because semirings are varieties and so any quotient by a semiring congruence is a semiring (see, for instance, [3]). To see that this quotient in fact defines a Kleene module, we must check for $a\in A$, $c\in C$, $m\in M$, and $n\in N$ that if $a(m,n)\leq(m,n)$ then $a^{*}(m,n)\leq(m,n)$, and that if $(m,n)c\leq(m,n)$ then $(m,n)c^{*}\leq(m,n)$. These inequalities follow immediately from the fact that $M$ is a left module over $A$ and $N$ is a right module over $M$, together with the fact that $(m,n)+(m,n)=(m,n)$.
@@
$\operatorname{Hom}(M\otimes_{B}N,P)$$\operatorname{Hom}(M,\operatorname{Hom}(N,P))$$\operatorname{Hom}(M^{\prime}\otimes_{B}N,P^{\prime})$$\operatorname{Hom}(M^{\prime},\operatorname{Hom}(N,P^{\prime}))$$\scriptstyle{\operatorname{Hom}(\alpha\otimes\text{id},\beta)}$$\scriptstyle{\operatorname{Hom}(\alpha,\operatorname{Hom}(N,\beta))}$
@@
$=\beta(\varphi(\alpha(m^{\prime})\otimes n)).$
@@
It remains to find an inverse to the natural transformation $\varphi\mapsto\overline{\varphi}$. The obvious candidate is given for $M\in{{}_{A}\text{Mod}_{B}}$, $P\in{{}_{B}\text{Mod}_{C}}$, $\psi\in\operatorname{Hom}(M,\operatorname{Hom}(N,P))$, $m\in M$, and $n\in N$ by $\hat{\psi}(m\otimes n)=\psi(m)(n)$. It is immediate that the transformation $\psi\mapsto\hat{\psi}$ is inverse to the transformation $\phi\mapsto\overline{\phi}$. It remains to verify that $\psi\mapsto\hat{\psi}$ is a natural transformation. For this we need to check that the following diagram commutes, where $M,M^{\prime}\in{{}_{A}\text{Mod}_{B}}$, $P,P^{\prime}\in{{}_{B}\text{Mod}_{C}}$, $\alpha\mathrel{:}M\rightarrow M^{\prime}$, $\beta\mathrel{:}P^{\prime}\rightarrow P$, and the horizontal arrows are given by application of the transformation $\psi\mapsto\hat{\psi}$.
@@
[…]
@@
The next results will be needed to verify that the Morita skeleton category from which we define Morita equivalence is indeed a category.
@@
Let $A$, $B$, $C$, $D$ be Kleene algebras, and ${{}_{A}}M_{B}$, ${{}_{B}}N_{C}$, ${{}_{C}}P_{D}$ be Kleene bimodules. Then
@@
For the isomorphism $({{}_{A}}M_{B}\otimes{{}_{B}}N_{C})\otimes{{}_{C}}P_{D}\cong{{}_{A}}M_{B}\otimes({{}_{B}}N_{C}\otimes{{}_{C}}P_{D})$ it suffices to check that the correspondence
$(m\otimes n)\otimes p\mapsto m\otimes(n\otimes p)$
for $m\in{{}_{A}}M_{B}$, $n\in{{}_{B}}N_{C}$, and $p\in{{}_{C}}P_{D}$; is well-defined, for then it extends linearly to a homomorphism which clearly has as inverse the homomorphism obtained from the (similarly well-defined) correspondence
$m\otimes(n\otimes p)\mapsto(m\otimes n)\otimes p.$
Suppose then that $m,m^{\prime}\in M$, $n,n^{\prime}\in N$, and $p,p^{\prime}\in P$, and that $(m\otimes n)\otimes p=(m^{\prime}\otimes n^{\prime})\otimes p^{\prime}$. This is witnessed by a sequence of steps of applying additivity in each variable and commuting scalars across the tensor product, that is to say, applying the generators of the congruence used to define the tensor product. Because scalars multiplying the middle variable may be pulled out of the inner tensor product (that is, $m^{\prime\prime}\otimes n^{\prime\prime}c=(m^{\prime\prime}\otimes n^{\prime\prime})c$ and $bn^{\prime\prime}\otimes p^{\prime\prime}=b(n^{\prime\prime}\otimes p^{\prime\prime})$ and then moved across the outer tensor product, the same sequence of operations can be applied to $m\otimes(n\otimes p)$ to obtain $m^{\prime}\otimes(n^{\prime}\otimes p^{\prime})$. This shows that the correspondence is well-defined and completes the proof of associativity of the tensor product up to isomorphism. The isomorphsims ${{}_{A}}A_{A}\otimes{{}_{A}}M_{B}\cong{{}_{A}}M_{B}$ and ${{}_{A}}M_{B}\otimes{{}_{B}}B_{B}\cong{{}_{A}}M_{B}$ are obtained by linearly extending the maps $1_{A}\otimes m\mapsto m$ and $m\otimes 1_{B}\mapsto m$ to homomorphisms and noting that these have respective inverses determined by $m\mapsto 1_{A}\otimes m$ and $m\mapsto 1_{B}\otimes B$. ∎
4 The Morita Skeleton of Semirings and of Kleene Algebras
The case of Morita equivalence for Kleene algebras is a straightforward adaptation of that for rings.
@@ 4 The Morita Skeleton of Semirings and of Kleene Algebras
The Morita skeleton of semirings is a $1$-category with objects isomorphism classes of semirings, [^5] Scott’s trick [17] may be used to represent these equivalence classes as sets. and for $A$ and $B$ semiringss the morphisms between them are all isomorphism classes of semiring $(A,B)$-bimodules (so this category is not locally small). The identity at an algebra $A$ is ${{}_{A}}A_{A}$, and composition of morphisms is given by the operation on isomorphism classes induced by tensor product (which is clearly well-defined because the tensor product is defined via an adjunction). Semirings $A$ and $B$ are called Morita-equivalent precisely when they are isomorphic in the Morita skeleton of semirings.
The Morita skeleton of Kleene algebras is the subcategory of the Morita skeleton of semirings with objects isomorphisms of Kleene algebras and morphisms isomorphism classes of Kleene bimodules.
In the case of rings one easily obtains a bicategory structure by taking $2$-cells to be homomorphisms between modules [2], and presumably this can be generalized to semirings and adapted to Kleene algebras, but since we do not use this extra structure we do not develop it.
The Morita skeleton (of semirings or of Kleene algebras) extends the usual category of semirings (Kleene algebras) and homomorphisms via the following notion of homomorphism module, which realizes homomorphisms as particular bimodules.
Definition 16.
Let $A$, $B$ be semirings (Kleene algebras) and $h\mathrel{:}A\rightarrow B$ be a homomorphism. The homomorphism module $E_{h}$ is a semiring (Kleene) $(A,B)$-bimodule whose elements and addition are given by the additive reduct of $B$, with the following multiplication actions on $b\in B$: For $a\in A$, $a\cdot_{E_{h}}b=h(a)\cdot_{B}b$. For $b^{\prime}\in B$, $b\cdot_{E_{h}}b^{\prime}=b\cdot_{B}b^{\prime}$.
The next proposition verifies that $E_{h}$ has the structure of a semiring (Kleene) bimodule.
@@
Let $A$ and $B$ be semirings (Kleene algebras) and $h:A\rightarrow B$ be a homomorphism. Then $E_{h}$ satisfies the axioms of a semiring (Kleene) $(A,B)$-bimodule.
Proof.
Clearly $E_{h}$ has the structure of a commutative idempotent monoid. The axioms for the right action of $B$ are immediate, and the axioms for the left action of $A$ follow because $h$ is a homomorphism. The fact that the two actions commute is an instance of associativity in $B$. In case $A$ and $B$ are Kleene algebras, the asterate inequality for the right action follows from the fact that $B$ is a Kleene algebra, while the asterate inequality on the left is verified by the following computation, where $a\in A$ and $b\in B$. If $a\cdot_{E_{h}}b=h(a)b\leq b$, then $h(a)^{*}b\leq b$ because $B$ is a Kleene algebra, and hence $a^{*}\cdot_{E_{h}}b=h(a^{*})b\leq b$ because $h(a)^{*}=h(a^{*})$ since $h$ is a homomorphism of Kleene algebras. ∎
Next we verify that the tensor product of homomorphism bimodules corrsponds in the Morita skeleton (of semirings or of Kleene algebras) to the homomorphism module of the composition, so that homomorphism modules do indeed give an extension of the category of semirings (Kleene algebras) and homomorphisms to the Morita skeleton.
Proposition 18.
Let $A$, $B$, $C$ be semirings (Kleene algebras) and $f\mathrel{:}A\rightarrow B$ and $g\mathrel{:}B\rightarrow C$ be homomorphisms. Then $E_{g\circ f}\cong E_{f}\otimes E_{g}$.
@@
$\phi(a\cdot_{E_{g\circ f}}c)$ $=1\otimes(g\circ f)(a)\cdot_{C}c$
@@
Now consider $b\otimes c\in E_{g}\otimes E_{f}$. Define $\psi\mathrel{:}E_{f}\otimes E_{g}\rightarrow E_{g\circ f}$ by $\psi(b\otimes c)=b\cdot{E_{g}}c$, and note that this is well-defined because $b\otimes c=1\otimes b\cdot_{E_{g}}c$ for $b\in B$, $c\in C$. It is clearly additive and commutes with the right action of $C$, so we demonstrate that it commutes with the left action of $A$. Let $a\in A$ and note that
@@
$=\psi(f(a)\otimes b\cdot_{E_{g}}c)$
$=f(a)(b\cdot_{E_{g}}c)$
$=a\psi(b\otimes c).$
@@
$\phi(\psi(b\otimes c))=\phi(b\cdot_{E_{g}}c)=1\otimes(b\cdot_{E_{g}}c)=b\otimes c$
@@
Another way of expressing Morita equivalence for rings is as equivalence of the corresponding module categories. This carries over to Kleene algebras, as we now argue. If $K$ is Morita-equivalent with $S$ via bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$, then for any left $K$-module ${{}_{K}}P$,
@@
As in the classical case, a Kleene algebra is Morita-equivalent with each of its finite-dimensional matrix rings. Because deterministic finite automata relative to a Kleene algebra $K$ correspond to square matrices over $K$, together with vectors of start and end states [11, Definition 12], this shows that reducts of automata in one Kleene algebra (forgetting the start and end states) correspond to single elements in a Morita-equivalent Kleene algebra. Unfortunately the author has not yet satisfactorily investigated concrete results which can be proved using this correspondence.
@@
The $(M,K)$-bimodule $K^{n}$ and its dual $K^{n\circ}$ will witness this equivalence. First consider $K^{n}\otimes K^{n\circ}$. For $1\leq i\leq n$ let $e_{i}$ be the “$i$th unit vector” of $K^{n}$, that is the tuple which is zero everywhere except position $i$, where it is $1$. Note that the set $\{e_{i}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n}$. Let $e_{i}^{\circ}$ be the dual of this module element, that is, $e_{i}^{\circ}(e_{i})=1$ and $e_{i}^{\circ}(e_{j})=0$ for $j\neq i$. It is easy to show that the set $\{e_{i}^{\circ}\mathrel{:}1\leq i\leq n\}$ freely generates $K^{n\circ}$. For $1\leq i,j\leq n$, let $E_{ij}\in M$ be the matrix which is zero everywhere except at position $(i,j)$, where it is $1$. Note that these freely generate $M$, and that the set $\{e_{i}\otimes e_{j}^{\circ}\mathrel{:}1\leq i,j\leq n\}$ freely generates $K^{n}\otimes K^{n\circ}$, because only elements of $K$ can be transported across the tensor product. Define a map $\phi\mathrel{:}K^{n}\otimes K^{n\circ}\rightarrow M$ by $\phi(e_{i}\otimes e_{j}^{\circ})=E_{ij}$; we show that $\phi$ is an isomorphism. Its inverse, naturally, will be given by $\psi(E_{ij})=e_{i}\otimes e_{j}^{\circ}$. By free generation $\phi$ and $\psi$ are well-defined homomorphisms, and it is immediate that they are inverses. Consequently $K^{n}\otimes K^{n\circ}\cong M$, which is half of a Morita equivalence.
For the other half we look at $K^{n\circ}\otimes K^{n}$, and our goal is to show this is isomorphic to $K$. For $e_{i}^{\circ}\otimes e_{j}\in K^{n\circ}\otimes K^{n}$, define $\alpha(e_{i}^{\circ}\otimes e_{j})$ to give the $ij$-entry of the identity matrix over $K$; that is, 0 if $i\neq j$ and $1$ if $i=j$. By free generation this determines a homomorphism $\alpha\mathrel{:}K^{n\circ}\otimes K^{n}\rightarrow K$. Define also $\beta\mathrel{:}K\rightarrow K^{n\circ}\otimes K^{n}$ by $\beta(a)=a(e_{1}^{\circ}\otimes e_{1})$ for $a\in A$. To complete the proof we show that $\alpha$ and $\beta$ are inverses. Certainly for $a\in A$ we have that
@@
To see that $\beta\circ\alpha=\text{id}$ it suffices to show that for any $i\leq n$, $e_{i}^{\circ}\otimes e_{i}=e_{1}^{\circ}\otimes e_{1}$ and for $j\leq n$, $j\neq i$, $e_{i}^{\circ}\otimes e_{j}=0$. For the latter let $E_{j}$ be the matrix obtained by replacing the $j$th row of the zero matrix by the “row vector” $e_{j}^{\circ}$. One computes easily that $E_{j}e_{j}=e_{j}$, and that $e_{i}^{\circ}E_{j}=0$. Hence
$e_{i}^{\circ}\otimes e_{j}=e_{i}^{\circ}\otimes E_{j}e_{j}=e_{i}^{\circ}E_{j}\otimes e_{j}=0\otimes e_{j}=0.$
For the case of $i=j$ we show that
@@
$\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$ $=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&0&0\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}=\begin{pmatrix}1&0&0\end{pmatrix}\otimes\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
$=\begin{pmatrix}1&0&0\end{pmatrix}\begin{pmatrix}1&1&1\\ 1&1&1\\ 1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}=\begin{pmatrix}1&1&1\end{pmatrix}\otimes\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$
@@
$e_{i}^{\circ}\otimes e_{i}$ $=e_{i}^{\circ}\overline{e}_{i}\otimes e_{i}=e_{i}^{\circ}\otimes\overline{e}_{i}e_{i}=e_{i}^{\circ}\otimes\sum_{i=1}^{n}e_{i}$
$=e_{i}^{\circ}\otimes\overline{1}\sum_{i=1}^{n}e_{i}=e_{i}^{\circ}\overline{1}\otimes\sum_{i=1}^{n}e_{i}=\left(\sum_{i=1}^{n}e_{i}^{\circ}\right)\otimes\sum_{i=1}^{n}e_{i}.$
@@
There is a well-known characterization of Morita equivalence for rings which extends to semirings [15, 9]. To state it we need the following definition:
@@
Let $R$ be a semiring and $e\in R$. The element $e$ is multiplicatively idempotent precisely when $e^{2}=e$, and full if and only if $ReR=R$.